# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ 4 Solution Methods
The main goal of this paper is to propose a new solution method for the Optimal and Fusion Problems under the A-criterion and the D-criterion based on the novel framework Boscia and assess its performance compared to several other convex MINLP approaches. In the following, we introduce the chosen MINLP solvers and state the necessary conditions and possible reformulations that are needed. We have chosen MINLP solution approaches which require relatively few changes to the formulations in (D-Opt) and (A-Opt). Hence, we are not investigating, for example, second-order cone formulations like in Sagnol & Harman (2015) in the scope of this paper.
Branch-and-Bound with Frank-Wolfe methods (Boscia).
The new framework introduced in (Hendrych et al., 2023) is implemented in the Julia package Boscia.jl. It is a Branch-and-Bound (BnB) framework that utilizes Frank-Wolfe methods to solve the relaxations at the node level. The Frank-Wolfe algorithm (Frank & Wolfe, 1956; Braun et al., 2022), also called Conditional Gradient algorithm (Levitin & Polyak, 1966), and its variants are first-order methods solving problems of the type:
@@
where $f$ is a convex, Lipschitz-smooth function and $\mathcal{X}$ is a compact convex set. These methods are especially useful if the linear minimization problem over $\mathcal{X}$ can be solved efficiently. The Frank-Wolfe methods used in Boscia are implemented in the Julia package FrankWolfe.jl, see (Besançon et al., 2022).
@@
At each iteration $t$, the Frank-Wolfe algorithm solves the linear minimization problem over $\mathcal{X}$ taking the current gradient as the linear objective, resulting in a vertex $v_{t}$ of $\mathcal{X}$. The next iterate $x_{t+1}$ is computed as a convex combination of the current iterate $x_{t}$ and the vertex $v_{t}$. Many Frank-Wolfe variants explicitly store the vertex decomposition of the iterate, henceforth called the active set. We utilize the active set representation to facilitate warm starts in Boscia by splitting the active set when branching.
@@
One novel aspect of Boscia is its use of a Bounded Mixed-Integer Linear Minimization Oracle (BLMO) as the Linear Minimization Oracle (LMO) in the Frank-Wolfe algorithm. The BLMO solves the mixed-integer linear problem over the feasible region $\mathcal{X}$ with additional bound constraints. Typically, the BLMO is a MIP solver but it can also be a combinatorial solver. This leads to more expensive node evaluations but has the benefit that feasible integer points are found from the root node and the feasible region is much tighter than the continuous relaxation for many problems. In addition, Frank-Wolfe methods can be lazified, i.e. calling the LMO at each iteration in the node evaluation can be avoided, see (Braun et al., 2017).
@@
In the case of OEDP, strong lazification is not necessary since the corresponding BLMO is very simple. The feasible region $\mathcal{P}$ is just the scaled probability simplex intersected with integer bounds, which is amenable to efficient linear optimization. Given a linear objective $d$, we first assign $\mathbf{x}=\mathbf{l}$ to ensure that the lower bound constraints are met. Next, we traverse the objective entries in increasing order, adding to the corresponding variable the value of $\max\{u-l,N-\sum(\mathbf{x})\}$. This way, we ensure that both the upper-bound constraints and the knapsack constraint are satisfied. The BLMO over the feasible set can also be cast as a simple network flow problem with $m$ input nodes connected to a single output node, which must receive a flow of $N$ while the edges respect the lower and upper bounds.
@@
Due to the convexity of the objective, the difference $\left\langle\nabla f(x_{t}),x_{t}-v_{t}\right\rangle$ is upper bounding the primal gap $f(x_{t})-f(x^{*})$ at each iteration. We call the quantity the dual gap (or the Frank-Wolfe gap). The dual gap can therefore be used as a stopping criterion. The error adaptiveness of the Frank-Wolfe algorithm can be exploited to a) solve nodes with smaller depth with a coarser precision and b) dynamically stop a node evaluation if the lower bound on the node solution exceeds the objective value at the incumbent.
@@
Observe that in contrast to the epigraph-based formulation approaches that generate many hyperplanes, our method works with the equivalent of a single supporting hyperplane given by the current gradient and moves this hyperplane until it achieves optimality. It is known that once the optimal solution is found, a single supporting hyperplane can be sufficient to prove optimality (as described e.g. for generalized Benders decomposition in Sahinidis & Grossmann (1991)). Finding this final hyperplane, however, may require adding many cuts beforehand at suboptimal iterations. In the case of the problems discussed in this paper, the constraint polytope is uni-modular. Adding hyperplanes created from the gradient will not maintain this structure and in consequence, yields a numerically more challenging MIP. Our approach, on the other hand, keeps the polytope and thereby its uni-modularity intact.
Outer Approximation (SCIP+OA).
@@
$\begin{split}\min_{t,\mathbf{x}}\;&t\\ \text{s.t. }&t\geq-\log\left(\phi_{p}(\mathbf{x})\right)\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$ (E-OEDP)
This approach approximates the feasible region of (E-OEDP) with linear cuts derived from the gradient of the non-linear constraints, in our case $\nabla f$. Note that this requires the information matrix $X(\mathbf{x})$ at the current iterate $\mathbf{x}$ to be positive definite, otherwise an evaluation of the gradient is not possible or rather it will evaluate to $\infty$. The implementation is done with the Julia wrapper of SCIP, (Bestuzheva et al., 2023). Observe that generating cuts that prohibit points leading to singular $X(\mathbf{x})$, we will refer to them as domain cuts, is non-trivial. Thus, this approach can only be used for the Fusion Problem where the corresponding information matrix is always positive definite.
LP/NLP Branch-and-Bound (Pajarito).
Another Outer Approximation approach, as implemented in Pajarito.jl (Coey et al., 2020), represents the non-linearities as conic constraints. This is particularly convenient in combination with the conic interior point solver Hypatia.jl (Coey et al., 2022c) as it implements the $\log\det$ cone (the epigraph of the perspective function of $\log\det$) directly:
$\mathcal{K}_{\log\det}:=\text{cl}\left\{(u,v,W)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\mathbb{S}_{++}^{n}\mid u\leq v\log\det(W/v)\right\}.$
@@
$\begin{split}\max_{t,\mathbf{x}}\;&t\\ \text{s.t. }&(t,1,X(x))\in\mathcal{K}_{\log\det}\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$
For the representation of the trace inverse, we utilize the dual of the separable spectral function cone (Coey et al., 2022b, Section 6):
$\mathcal{K}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathcal{Q})\mid u\geq v\varphi(w/v)\right\}.$
For our purposes, $\mathcal{Q}$ is the PSD cone and the spectral function $\varphi$ is the negative square root whose convex conjugate is precisely the trace inverse, see (Coey et al., 2022a, Table 1) [^2] For further details, see this discussion Kapelevich (2023).:
$\mathcal{K}^{*}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathbb{S}_{+}^{n})\mid v\geq u/4\operatorname{Tr}((w/u)^{-1})\right\}.$
The conic formulation of (A-Opt) is therefore
@@
Additionally, the conic formulation allows for the computation of domain cuts for $\mathbf{x}$. Hence, this solver can be used on all problems. Note that we use HiGHS (Huangfu & Hall, 2018) as a MIP solver within Pajarito.
@@
The most general solver strategy for (OEDP) with matrix means criteria was introduced in Ahipaşaoğlu (2021). Like Boscia , it is a Branch-and-Bound-based approach with a first-order method to solve the node problems. The first-order method in question is a coordinate-descent-like algorithm. We refer to this approach as Co-BnB in the rest of this paper. As the termination criterion, this method exploits the fact that the objective function is a matrix mean and shows the connection of the resulting optimization problem:
@@
to the generalization of the Minimum Volume Enclosing Ellipsoid Problem (MVEP) (Ahipaşaoğlu, 2021). The variables $w$ can be interpreted as a probability distribution and the number of times the experiments are to be run is denoted by $wN$. Concerning (OEDP), one can say $x=wN$. In the case of the Fusion Problem, the already completed experiments have to be explicitly added as variables with fixed values.
Note that we have improved and adapted the step size rules within the first-order method, see Appendix 0.B. Further, note that the solver was developed for instances with a plethora of experiments and very few parameters. The solver employs the simplest Branch-and-Bound strategy, i.e. utilizing the most fractional branching rule and traversing the tree using the minimum lower bound. In the next section, we will see that the method works well in cases where $n$ is small but struggles if $n$ increases.
@@ 5 Computational Experiments
5.0.1 Experimental Setup
@@
For the instance generation, we choose the number of experiments $m\in\{50,60,80,100,120\}$, the number of parameters $n\in\{\lfloor m/4\rfloor,\lfloor m/10\rfloor\}$, and the number of allowed experiments $N=\lfloor 1.5n\rfloor$ for the Optimal Problems and $N\in[m/20,m/3]$ for the Fusion Problems. The lower bounds are zero. Note that for the Fusion Problems, the fixed experiments are encoded in a separate matrix. The upper bounds are randomly sampled between 1 and $N/3$ for (A-Opt) and (D-Opt). In the Fusion case, they are sampled between 1 and $m/10$. We generate both independent and correlated data. Also, note that the matrices generated are dense. The number of random seeds is 5. In total, there are 50 instances for each combination of problem and data set.
@@
Experiments were run on a cluster equipped with Intel Xeon Gold 6338 CPUs running at 2 GHz and a one-hour time limit. The source code is hosted on GitHub [^3] https://github.com/ZIB-IOL/OptimalDesignWithBoscia.
Start Solution
@@
Note that both the objectives (D-Opt) and (A-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank. This is the case for the Fusion Problem, not necessarily for the Optimal Problem. Both Boscia and Co-BnB require a feasible starting point $\mathbf{z}_{0}$. For its construction, we find a set $S\subset[m]$ of $n$ linearly independent experiments, i.e. $n$ linearly independent rows of $A$. Assign those experiments their upper bound. If the sum $\left\langle 1,z_{0}\right\rangle$ exceeds $N$, remove 1 from the experiment with the largest entry. If the sum is less than $N$, pick an experiment in $[m]\backslash S$ at random and assign it as many runs as possible. Repeat until the sum $\left\langle 1,z_{0}\right\rangle$ is equal to $N$. Note that due to the monotonic progress of both first-order methods, the current iterate will never become domain infeasible, i.e. $X(\mathbf{x})$ will not become singular. If the iterate is domain infeasible after branching, we discard that node.
5.0.2 Results
An overview of the results of the computational experiments is given in Table 1. The new framework Boscia solves the most instances by far. In comparison to the Outer Approximation methods, it terminates nearly twice as often. The Co-BnB fares better. Nonetheless, it solves fewer instances to optimality than Boscia. The notable exception is the A-Fusion Problem with correlated data. Note that in general, it fares well for the instances where $n=\lfloor m/10\rfloor$ as it was designed for such problems. It struggles for the instances where $n=\lfloor m/4\rfloor$.
In terms of time, Boscia also shows promising results, especially for instances of larger scale. For small-scale instances, the Outer Approximation approaches are fast, in some cases faster than Boscia. The graphical view of the number of solved instances over time is shown in Figure 1 for the Optimal Problems and in Figure 2 for the Fusion Problems.
(a) A-Optimal Problem
(b) D-Optimal Problem

### v6

@@ 4 Solution methods
The main goal of this paper is to propose a new solution method for the Optimal and Fusion Problems under the A-criterion and D-criterion based on the novel framework Boscia.jl and assess its performance compared to several other convex MINLP approaches. In the following, we introduce the chosen MINLP solvers and state the necessary conditions and possible reformulations that are needed.
Branch-and-bound with Frank-Wolfe algorithms (Boscia).
The new framework introduced in Hendrych et al. (2025b) is implemented in the Julia package Boscia.jl. It is a Branch-and-Bound (BnB) framework that utilizes Frank-Wolfe methods to solve the relaxations at the node level. The Frank-Wolfe algorithm (Frank & Wolfe, 1956; Braun et al., 2022), also called Conditional Gradient algorithm (Levitin & Polyak, 1966), and its variants are first-order methods solving problems of the type:
@@
where $f$ is a convex, Lipschitz-smooth function and $\mathcal{X}$ is a compact convex set. These methods are especially useful if linear minimization problems over $\mathcal{X}$ can be solved efficiently. The Frank-Wolfe methods used in Boscia.jl are implemented in the Julia package FrankWolfe.jl, see Besançon et al. (2022; 2025).
@@
At each iteration $t$, Frank-Wolfe solves the linear minimization problem over $\mathcal{X}$ taking the current gradient as the linear objective, resulting in a vertex $v_{t}$ of $\mathcal{X}$. The next iterate $x_{t+1}$ is computed as a convex combination of the current iterate $x_{t}$ and the vertex $v_{t}$. Many Frank-Wolfe variants explicitly store the vertex decomposition of the iterate, henceforth called the active set. We utilize the active set representation to facilitate warm starts in Boscia by splitting the active set when branching.
@@
One novel aspect of Boscia.jl is its use of a Bounded Mixed-Integer Linear Minimization Oracle (BLMO) as the Linear Minimization Oracle (LMO) in Frank-Wolfe. Typically, the BLMO is a MIP solver but it can also be a combinatorial solver. This leads to more expensive node evaluations but has the benefit that feasible integer points are found from the root node and the feasible region is much tighter than the continuous relaxation for many problems. In addition, Frank-Wolfe methods can be lazified, i.e. calling the LMO at each iteration in the node evaluation can be avoided, see (Braun et al., 2017).
@@
In the case of OEDP, strong lazification is not necessary since the corresponding BLMO is very simple. The feasible region $\mathcal{P}$ is just the scaled probability simplex intersected with integer bounds, which is amenable to efficient linear optimization. Given a linear objective $d$, we first assign $\mathbf{x}=\mathbf{l}$ to ensure that the lower bound constraints are met. Next, we traverse the objective entries in increasing order, adding to the corresponding variable the value of $\max\{u_{i}-l_{i},N-\sum(\mathbf{x})\}$. This way, we ensure that both the upper-bound constraints and the knapsack constraint are satisfied. The LMO over the feasible set can also be cast as a simple network flow problem with $m$ input nodes connected to a single output node, which must receive a flow of $N$ while the edges respect the lower and upper bounds. The scaled probability simplex also allows for a simple heuristic taking a fractional point and outputting an integer feasible point on the probability simplex.
@@
Due to the convexity of the objective, the difference $\left\langle\nabla f,x_{t}-v_{t}\right\rangle$ is upper bounding the primal gap $f(x_{t})-f(x^{*})$ at each iteration. We call the quantity the dual gap (or the Frank-Wolfe gap). The dual gap can therefore be used as a stopping criterion. Frank-Wolfe’s error adaptiveness can be exploited to a) solve nodes with smaller depth with a coarser precision and b) dynamically stop a node evaluation if the lower bound on the node solution exceeds the incumbent.
@@
Observe that in contrast to the epigraph-based formulation approaches that generate many hyperplanes, this method works with the equivalent of a single supporting hyperplane given by the current gradient and moves this hyperplane until it achieves optimality. It is known that once the optimal solution is found, a single supporting hyperplane can be sufficient to prove optimality (described e.g. for generalized Benders in Sahinidis & Grossmann (1991)). Finding this final hyperplane, however, may require adding many hyperplanes beforehand at suboptimal iterations. In the case of the problems discussed in this paper, the constraint polytope is uni-modular. Adding hyperplanes created from the gradient will not maintain this structure and in consequence, yields a numerically more challenging MIP. Our approach, on the other hand, keeps the polytope and thereby its uni-modularity intact.
A new development during this work was the implementation of a domain oracle in Boscia.jl. As stated earlier, not all points in the feasible region are domain feasible for the objective functions. Thus, we cannot start Boscia (and Frank-Wolfe) at a random point. We therefore added the option of a domain oracle and a domain point function. The former checks whether a given point is domain feasible, in our case by checking if the associated information matrix is regular. The latter provides a domain feasible point given the feasible region and additional bound constraints from the node. The returned point has to respect the global constraints as well as the node specific constraints. The domain point function is called during branching if the splitting of the active set results in an invalid point. A small projection problem is then solved, also using Frank-Wolfe, to find a new point within the intersection of the feasible region and the domain. Note that we stop the Frank-Wolfe method once we are in the intersection.
For the line search problem within Frank-Wolfe, we utilize the Secant line search which has shown a superior performance for generalized self concordant functions, see Hendrych et al. (2025a).
Outer approximation (SCIP OA).
@@
$\begin{split}\min_{t,\mathbf{x}}\;&t\\ \text{s.t. }&t\geq\log\left(\phi_{p}(\mathbf{x})\right)\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$ (E-OEDP)
This approach approximates the feasible region of (E-OEDP) with linear cuts derived from the gradient of the non-linear constraints, in our case $\nabla f$. Note that this requires the information matrix $X(\mathbf{x})$ at the current iterate $\mathbf{x}$ to be positive definite, otherwise an evaluation of the gradient is not possible or rather it will evaluate to $\infty$. The implementation is done with the Julia wrapper of SCIP, (Bestuzheva et al., 2021; 2023). Note that generating cuts that prohibit points leading to singular $X(\mathbf{x})$, we will refer to them as domain cuts, is non-trivial. Thus, we are applying this approach to the Fusion Problems only since their information matrix is always positive definite.
LP/NLP branch-and-bound with direct conic formulation (Direct Conic).
Another Outer Approximation approach, as implemented in Pajarito.jl (Coey et al., 2020), represents the non-linearities as conic constraints. This is particularly convenient in combination with the conic interior point solver Hypatia.jl (Coey et al., 2022b) as it implements the $\log\det$ cone (the epigraph of the perspective function of $\log\det$) directly.
$\mathcal{K}_{\log\det}:=\text{cl}\left\{(u,v,W)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\mathbb{S}_{++}^{n}\mid u\leq v\log\det(W/v)\right\}$
@@
$\begin{split}\max_{t,\mathbf{x}}\;&t\\ \text{s.t. }&(t,1,X(\mathbf{x}))\in\mathcal{K}_{\log\det}\\ &\mathbf{x}\in\mathcal{P}\\ &t\in{\mathbb{R}},\mathbf{x}\in{\mathbb{Z}}^{m}.\end{split}$
For the representation of the trace inverse, we utilize the dual of the separable spectral function cone (Coey et al., 2022a, Section 6):
$\mathcal{K}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathcal{Q})\mid u\geq v\varphi(w/v)\right\}$
For our purposes, $\mathcal{Q}$ is the PSD cone and the spectral function $\varphi$ is the negative square root whose convex conjugate is precisely the trace inverse, see (Coey et al., 2023, Table 1) and (Pilanci et al., 2015, Collary 1) [^4] For further details, see this discussion Kapelevich (2023)..
$\mathcal{K}^{*}_{\text{sepspec}}:=\text{cl}\left\{(u,v,w)\in{\mathbb{R}}\times{\mathbb{R}}_{>0}\times\text{int}(\mathbb{S}_{+}^{n})\mid v\geq u/4\operatorname{Tr}((w/u)^{-1})\right\}$
The conic formulation of the (A-Opt) is therefore
@@
Additionally, the conic formulation allows the computation of domain cuts for $\mathbf{x}$. Hence, this solver can be used on all problems. Note that we use HiGHS (Huangfu & Hall, 2018) as a MIP solver within Pajarito.jl.
LP/NLP branch-and-bound with second-order cone formulation (SOCP).
The conic solver from the previous approach was developed fairly recently. A previous conic approach was introduced first in Sagnol (2011) for the continuous case and later extended to the mixed-integer case in Sagnol & Harman (2015). They showed OEDP under the A-criterion and D-criterion can be formulated as a second-order cone program (SOCP). The SOCP model for (D-Opt) is given by
$\begin{split}\max_{\mathbf{w},Z,T,J}\;&\prod\limits_{j=1}^{n}\left(J_{jj}\right)^{\frac{1}{n}}\\ \text{s.t. }&\sum\limits_{i=1}^{m}A_{i}Z_{i}=J\\ &J\text{ is a lower triangle matrix}\\ &\left\|Z_{i}\mathbf{e}_{j}\right\|^{2}\leq T_{ij}\mathbf{w}_{i}\quad\forall i\in[m],j\in[n]\\ &\sum\limits_{i=1}^{m}T_{ij}\leq J{jj}\quad\forall j\in[n]\\ &T_{ij}\geq 0\quad\forall i\in[m],j\in[n]\\ &\sum\limits_{i=1}^{m}\mathbf{w}_{i}=N,\;\mathbf{w}\geq 0\end{split}$
where $J$ is a $n\times n$ matrix, $Z$ and $T$ are $m\times n$ matrices. Observe that the problem size has increased significantly with both the number of variables and number of contraints being $\geq m(2n+1)$.
The SOCP formulation for (A-Opt) is
$\begin{split}\min_{\mathbf{w},\bm{\mu},Y}\;&\sum\limits_{i=1}^{m}\mu_{i}\\ \text{s.t. }&\sum\limits_{i=1}^{m}A_{i}Y_{i}=\sum\limits_{i=1}^{m}\bm{\mu}_{i}\\ &\left\|Y_{j}\right\|^{2}\leq\bm{\mu}_{j}\mathbf{w}_{i}\quad\forall i\in[m]\\ &\bm{\mu}\geq 0,\;\mathbf{w}\geq 0\\ &\sum\limits_{i=1}^{m}\mathbf{w}_{i}=N\end{split}$
where $Y$ is a $m\times n$ matrix and $\bm{\mu}$ is a $m\times 1$ vector. The problem size does not increase as much, with $m(2+n)$ variables and $23m+$ 2 constraints.
To solve the SOCP formulation, we again utilize Pajarito.jl with HiGHS as a MIP solver and Hypatia.jl as a conic solver. For the terms involving the norm, we make use of Hypatia.jl’s EpiNormEuclCone and EpiPerSquareCone cones which do not require us to deconstruct the norm terms into second order cones.
@@
The most general solver strategy for (OEDP) with matrix means criteria was introduced in Ahipaşaoğlu (2021). Like Boscia.jl, it is a Branch-and-Bound-based approach with a first-order method to solve the node problems. The first-order method in question is a coordinate-descent-like algorithm. As the termination criterion, this method exploits that the objective function is a matrix mean and shows the connection of the resulting optimization problem
@@
to the generalization of the Minimum Volume Enclosing Ellipsoid Problem (MVEP) (Ahipaşaoğlu, 2021). The variables $w$ can be interpreted as a probability distribution and the number of times the experiments are to be run is $wN$. Concerning (OEDP), one can say $x=wN$.
Note first that we have reimplemented the method in Julia and modified it so that we solve the formulation as depicted in (OEDP). Secondly, we have improved and adapted the step size rules within the first-order method, see Section B. Note further that the solver was developed for instances with a plethora of experiments and very few parameters. The solver employs the simplest Branch-and-Bound tree, i.e. with the most fractional branching rule and traverses the tree with respect to the minimum lower bound. In the next section, we will see that the method works well in cases where $n$ is small but struggles if the value of $n$ increases.
@@ 5 Computational experiments
Furthermore, we run the Optimal and Fusion Problems under the GTI-criterion for different values of $p$ using permissible solvers. Lastly, we showcase the performance of Boscia.jl and Co-BnB on problems of large dimension, i.e. number of variables between 300 and 500.
Experimental setup.
@@
For the instance generation, we choose the number of experiments $m\in\{50,60,80,\linebreak 100,120\}$, the number of parameters $n\in\{\lfloor m/4\rfloor,\lfloor m/10\rfloor\}$, and the number of allowed experiments $N=\lfloor 1.5n\rfloor$ for the Optimal Problems and $N\in[m/20,m/3]$ for the Fusion Problems. The lower bounds are zero. Note that for the Fusion Problems, the fixed experiments are encoded in a separate matrix. The upper bounds are randomly sampled between 1 and $N/3$ for (A-Opt) and (D-Opt). In the Fusion case, they are sampled between 1 and $m/10$. We generate both independent and correlated experiment data. Also, note that the matrices generated are dense. Seeds used are $[1,2,3,4,5]$. In total, there are 50 instances for each combination of problem and data set.
@@
Experiments were run on a cluster equipped with Intel Xeon Gold 6338 CPUs running at 2 GHz and a one-hour time limit. The Julia version used is 1.10.2, Boscia.jl v0.1.35, FrankWolfe.jl v0.4.13, HiGHS.jl v1.15.0, Hypatia.jl v0.7.4, Pajarito.jl v0.8.2, SCIP.jl v0.12.3. Note that we use the unregistered package PajaritoExtras.jl to model the dual of the separable spectral function cone. The source code is hosted on GitHub [^5] https://github.com/ZIB-IOL/OptimalDesignWithBoscia.
Start solution.
@@
Note that both the objectives (D-Opt) and (A-Opt) are only well defined if the information matrix $X(\mathbf{x})$ has full rank. This is the case for the Fusion Problem, not necessarily for the Optimal Problem. Both Boscia.jl and Co-BnB require a feasible starting point $\mathbf{z}_{0}$. For its construction, we find a set $S\subset[m]$ of $n$ linearly independent experiments, i.e. $n$ linearly independent rows of $A$. Assign those experiments their upper bound. If the sum $\sum\mathbf{z}_{0}$ exceeds $N$, remove 1 from the experiment with the largest entry. If the sum is less than $N$, pick an experiment in $[m]\backslash S$ at random and assign it as many runs as possible. Repeat until the sum is equal to $N$. Note that due to the monotonic progress of both first-order methods, the current iterate will never become domain infeasible, i.e. singular.
Results.
An overview of the results of the computational experiments is given in Table 2 and extensive results can be found in Section D. The new framework Boscia.jl shows a superior performance compared to the other methods both in terms of time and number of instances solved to optimality. In comparison to the Outer Approximation methods, for example, it solves nearly twice as many instances.
Co-BnB is the only competitive method. It mostly solves fewer instances than Boscia.jl, the exception being the A-Fusion Problem with both data sets. With independent data, Co-BnB is faster solving the same number of instances. In case of the correlated data set, Co-BnB solves roughly twice as many instances as Boscia.jl. In Table 5, we see that Boscia.jl and Co-BnB have similar magnitudes in number of nodes for both A-Fusion Problems whereas for all other problems Boscia.jl has significantly fewer nodes than Co-BnB. As stated previously, node evaluation in Boscia.jl is more computationally expensive. The fact that the magnitudes of nodes is similar points to the optimal solution being strictly in the interior. Note that in general, Co-BnB fares well for the instances where $n=\lfloor m/10\rfloor$ as it was designed with such problems. It struggles for the instances where $n=\lfloor m/4\rfloor$. In Figure 3, we see that the absolute gap of Co-BnB increasing drastically as the fraction between number of parameters and number experiments grow larger. While Boscia.jl also shows an increase, it is much more moderate. Co-BnB implements no advanced Branch-and-Bound tree specifications like a better traverse strategy or branching strategy. A greater value of $N$ naturally increases the size of the tree and the number of nodes to be processed. Boscia.jl has the advantage here since it finds many integer feasible points while solving the relaxations which have the potential to improve the incumbent. A better incumbent, in turn, lets us prune non-improving nodes early on. Overall, fewer nodes have to be investigated by Boscia.jl.
(a) A-Optimal Problem with independent data
(b) D-Optimal Problem with independent data
(c) A-Optimal Problem with correlated data
(d) D-Optimal Problem with correlated data
Figure 3: Geometric mean and geometric standard deviation of the absolute gap between the incumbent and lower bound over the conditioning of the problem.
In terms of time, Boscia.jl also shows promising results, especially for instances of larger scale. For small-scale instances, the Outer Approximation approaches are fast, in some cases faster than Boscia.jl. This can be observed in Figures 4 and 5. The plots also highlight that the Outer Approximation Schemes solve fewer instances than the two Branch-and-Bound Approaches. It should be noted that SCIP + OA simply timed out often. On the other hand, the MIP solvers for the conic formulations with Pajarito.jl often struggled to solve the resulting MIP problems. SOCP, for example, only solved about a quarter of the instances of the D-Fusion Problem and D-Optimal Problem. In addition, the many solutions provided by SOCP in case of a timeout were not always domain feasible, and thus, the original objective function could not be evaluated at the provided solutions. The superiority of the Branch-and-Bound approaches stems from the fact that the relaxations of both of them keep the simple feasible region intact. Whereas the Outer Approximation schemes add many additional constraints, i.e. cuts, resulting in larger LPs to be solved, see Table 5 for the average number of cuts. Furthermore, these cuts are dense, i.e. have many nonzero entries, leading to an increased difficulty for the LP solvers.
(a) A-Optimal Problem with independent data
(b) D-Optimal Problem with independent data
(c) A-Optimal Problem with correlated data
(d) D-Optimal Problem with correlated data